### v1

# From Misleading Queries to Accurate Answers: A Three-Stage Fine-Tuning Method for LLMs

Abstract
Large language models (LLMs) exhibit excellent performance in natural language processing (NLP), but remain highly sensitive to the quality of input queries, especially when these queries contain misleading or inaccurate information. Existing methods focus on correcting the output, but they often overlook the potential of improving the ability of LLMs to detect and correct misleading content in the input itself. In this paper, we propose a novel three-stage fine-tuning method that enhances the ability of LLMs to detect and correct misleading information in the input, further improving response accuracy and reducing hallucinations. Specifically, the three stages include (1) training LLMs to identify misleading information, (2) training LLMs to correct the misleading information using built-in or external knowledge, and (3) training LLMs to generate accurate answers based on the corrected queries. To evaluate our method, we conducted experiments on three datasets for the hallucination detection task and the question answering (QA) task, as well as two datasets containing misleading information that we constructed. The experimental results demonstrate that our method significantly improves the accuracy and factuality of LLM responses, while also enhancing the ability to detect hallucinations and reducing the generation of hallucinations in the output, particularly when the query contains misleading information. We will publicly release our code upon acceptance.
From Misleading Queries to Accurate Answers:
A Three-Stage Fine-Tuning Method for LLMs
Guocong Li1, Weize Liu1, Yihang Wu1, Ping Wang2, Shuaihan Huang1, Hongxia Xu1,3 Jian Wu1 1Zhejiang University    2Renmin University of China 3AI Research Center, WeDoctor Cloud
Figure 1: When LLMs directly answer a query containing misleading information (Left side), they produce incorrect responses. However, when LLMs identify and correct the misleading information in the query before generating a response, they correctly answer it  (Right side). The misleading information is marked in red, while the corresponding corrections are marked in blue.
1 Introduction
LLMs have achieved significant success in the field of NLP, particularly excelling in NLG, where they are capable of generating coherent and reasonable responses based on user queries (Li et al., 2024b; Xuanfan and Piji, 2023). Although LLMs have made significant progress in generation ability, they are highly sensitive to the quality of the input, and the quality of the generated responses is highly susceptible (Pan et al., 2023; Wu et al., 2024a). Users often provide LLMs with queries that are imprecise or may even contain various forms of misleading information (Nie et al., 2020). Misleading information refers to inaccurate content within the input that can mislead the model during its reasoning and generation processes. When such information is present in the input, the likelihood of the model generating incorrect responses or hallucinations increases significantly (Yu et al., 2024; Song et al., 2024). This challenge is particularly prominent in knowledge-intensive domains such as medicine, law, and education, where inputs often contain inaccurate information, and the accuracy of responses is critically important (Barnard et al., 2023). Therefore, mitigating the impact of misleading information in the input on the performance of LLMs is an urgent and important issue that needs to be addressed.
Current training methods for LLMs primarily focus on correcting the output after generation. An important research direction is the use of the retrieval-augmented generation (RAG) framework, which leverages external knowledge bases to verify and improve the output (Lewis et al., 2020; Gao et al., 2023). However, RAG typically requires retrieval from large external knowledge bases, a process that can be time-consuming and may not meet the low-latency requirements, especially in real-time application scenarios (Li et al., 2024a; Jin et al., 2024). In addition, the self-correction mechanism has also garnered attention, where LLMs iteratively generate and self-evaluate to critique and correct their own outputs (Madaan et al., 2024). This method has the potential to improve the accuracy and quality of the results, but it also has the drawback that the effectiveness of the correction depends on the model’s accuracy in evaluating the results and may require external feedback for assistance (Kamoi et al., 2024). These existing methods focus on correcting the content generated by LLMs, without considering the correction of misleading information in the input. Although the presence of misleading information in the input may lead to incorrect responses, our proposed method is capable of identifying and detecting errors in the query, subsequently generating the correct answer, as shown in Figure 1. However, no existing work has considered improving the performance of LLMs by enhancing their ability to identify and correct misleading questions.
Therefore, we propose a novel three-stage fine-tuning method that enhances LLMs’ ability to detect and correct misleading information in the input. This method enhances the accuracy and robustness of LLMs in processing inputs containing misleading information, while mitigating the negative impact of misleading information on model outputs.
The fine-tuning method that we propose consists of the following three training stages:
1) Query detection training: Training LLMs to determine whether the input query contains misleading information based on built-in knowledge or external knowledge;
2) Query correction training: Training LLMs to combine built-in knowledge or external knowledge to correct the misleading information in the query and generate the correct query;
3) Query answering training: Training LLMs to generate accurate and reliable answers based on the corrected query.
To evaluate the performance of LLMs in answering queries containing misleading information, we construct two misleading datasets, HaluEval-QAmis and CQAmis, based on the QA subset of the HaluEval dataset and the CQA dataset, respectively. Further, We conduct experiments on the misleading datasets, as well as on the three original datasets for the hallucination detection task and the question answering task. The experimental results show that when input contains misleading information, the performance of the trained model using our method is much higher than that of the original model, demonstrating that our method effectively reduces the impact of misleading information in the input on model performance. Furthermore, regardless of whether the input contains misleading information, our method significantly improves the accuracy and factuality of LLM responses, as well as the ability to detect hallucinations in the responses, while effectively mitigating the hallucinations of the model.
2 A Fine-Tuning Method to Mitigate the Impact of Misleading Information
Figure 2: The framework diagram of our method. In the first stage, we train LLMs to determine whether the query contains misleading information. In the second stage, for queries containing misleading information, we train LLMs to correct the misleading information in the query to obtain a correct query. In the third stage, we train LLMs to generate the correct answer based on the corrected query or the original query without misleading information. Both the Correct Query and Correct Answer can be directly derived from the content of the original dataset.
The framework of the proposed method is illustrated in Figure 2. It involves training LLMs to sequentially perform the following tasks: detecting misleading information in the input query, correcting the misleading query, and generating accurate answers based on the corrected query, thereby improving the model’s accuracy and robustness when handling inputs containing misleading information.
2.1 Construction of Misleading Query Datasets
We used the Qwen2.5-72B-Instruct model to generate a dataset containing misleading information based on the original queries. Specifically, we constructed the misleading query dataset HaluEval-QAmis by augmenting the queries from the QA subset of the HaluEval dataset with misleading content, leveraging the knowledge from the original queries in the HaluEval dataset. Similarly, based on the CQA dataset, we used the Qwen2.5-72B-Instruct model to modify each question in the original dataset to contain misleading information, thereby constructing the misleading query dataset CQAmis. For each dataset, we designed corresponding prompts to ensure that the newly generated misleading query datasets introduced misleading information while preserving the naturalness and complexity of the original queries. The prompt templates and data examples used are provided in Appendix B. At the same time, to ensure that the generated misleading query datasets are of higher quality, we adopted a batch generation strategy, generating three misleading queries per batch, and filtered the generated data based on the following two criteria:
Sentence Similarity
The similarity between the generated misleading query and the original query is measured using edit distance, which is defined as the minimum number of operations (including insertion, deletion, and substitution) required to transform one string into another (Niu et al., 2018). Let the original query be $q_{\text{ori}}$ and the generated misleading query be $q_{\text{mis}}$. The sentence similarity $S_{\text{sim}}$ is defined as:
$S_{\text{sim}}=1-\frac{\text{EditDistance}(q_{\text{ori}},q_{\text{mis}})}{\max(\text{Length}(q_{\text{ori}}),\text{Length}(q_{\text{mis}}))},$ (1)
where $\text{EditDistance}(q_{\text{ori}},q_{\text{mis}})$ represents the edit distance between $q_{\text{ori}}$ and $q_{\text{mis}}$, and Length denotes the length of the sentence. We retain misleading queries with surface similarity $S_{\text{sim}}$ higher than the threshold $r_{\text{sim}}$ to ensure that the generated queries are close to the original queries.
Answer Error Rate
The answer error rate measures whether the model generates incorrect answers when responding to generated misleading queries. We input the query into the model $N$ times, record the number of incorrect answers $N_{\text{error}}$, and define the error rate as:
$E_{\text{error}}=\frac{N_{\text{error}}}{N},$ (2)
where $E_{\text{error}}$ denotes the error rate. We only retain the misleading queries with an answer error rate $E_{\text{error}}$ above the threshold $r_{\text{error}}$ to ensure that these queries have a high level of misleading content.
Data Filtering Strategy
For each query, we randomly generate three misleading queries that satisfy $S_{\text{sim}}$ > 0.8 and $E_{\text{error}}$ > 0.5, and then retain the one with the highest value of $S_{\text{sim}}+E_{\text{error}}$. An example of randomly generated misleading queries is presented in Appendix  C. This ensures that the generated misleading queries are both highly similar to the original queries and contain misleading information, thereby enabling more effective training and evaluation of LLMs’ performance when faced with highly deceptive and subtle misleading queries.
2.2 Query Detection Training
Since misleading information in the query can significantly affect the quality of LLMs’ responses, the goal of this stage is to train the model to identify whether misleading information exists in the given query.
Input
A query. (If the dataset contains background knowledge, the background knowledge is also input to the model, such as HaluEval dataset.)
Output Objective
YES (indicating that the query contains misleading information) or NO (indicating that the query contains no misleading information).
Training Objective
The training objective is to minimize the cross-entropy loss function:
$\begin{split}\mathcal{L}_{\text{check}}=&-\frac{1}{N}\sum_{i=1}^{N}\left(y_{\text{true},i}\log(y_{\text{pred},i})\right.\\ &\quad+\left.(1-y_{\text{true},i})\log(1-y_{\text{pred},i})\right)\end{split},$ (3)
where $N$ is the number of samples in each batch, $y_{\text{true},i}$ is the true label of sample $i$, with a value of 1 (indicating “YES”) or 0 (indicating “NO”), and $y_{\text{pred},i}$ is the model’s predicted probability that the $i$-th sample is labeled as “YES”.
2.3 Query Correction Training
Queries labeled as containing misleading information (“YES”) in the query detection stage will proceed to the query correction stage. The goal of this stage is to correct the queries by leveraging the knowledge provided by the dataset or the model’s internal knowledge, transforming them into versions that do not contain misleading information, thereby improving the query quality.
Input
A misleading query. (If the dataset contains background knowledge, the background knowledge is also input to the model, such as HaluEval dataset.)
Output Objective
Corrected Query.
Training Objective
The loss function in the query correction stage uses semantic similarity rather than edit distance, as misleading information primarily affects semantic similarity rather than edit distance (Wu et al., 2022). This approach is used to train the model to generate high-quality corrected queries that are semantically consistent with the original correct queries. In this study, we use Sentence-BERT to obtain semantic vector representations (Reimers, 2019). Furthermore, we ensure that corrected queries are semantically and logically consistent with correct queries by minimizing the distance between corrected and correct queries in the semantic space. Mathematically, the loss function is formulated as:
$\mathcal{L}_{\text{correction}}=-\frac{1}{N}\sum_{i=1}^{N}\left(1-\frac{\mathbf{u}_{i}\cdot\mathbf{v}_{i}}{\|\mathbf{u}_{i}\|\|\mathbf{v}_{i}\|}\right),$ (4)
where $N$ is the number of training samples in each batch, and $\mathbf{u}_{i}$ and $\mathbf{v}_{i}$ are the semantic vector representations of the $i$-th correct query and corrected query, respectively.
2.4 Query Answering Training
The goal of this stage is to generate accurate and reliable answers based on the corrected query or the original query.
Input
The corrected query or the original query.
Output Objective
Correct Answers.
Training Objective
The model-generated answers should be aligned with the ground truth answers. The training objective is to minimize the cross-entropy loss function. Given the input sequence of the $i$-th data sample as $X^{(i)}=x_{1}^{(i)},x_{2}^{(i)},\dots,x_{T_{i}}^{(i)}$, the loss function is defined as:
$\begin{split}&\mathcal{L}_{\text{generation}}=\\ &-\frac{1}{N}\sum_{i=1}^{N}\left(\frac{1}{T_{i}}\sum_{t=1}^{T_{i}}\log P(y_{t}^{(i)}|x_{1}^{(i)},x_{2}^{(i)},\dots,x_{t-1}^{(i)})\right)\end{split},$ (5)
where $N$ is the number of training samples in each batch, $P(y_{t}^{(i)}|x_{1}^{(i)},x_{2}^{(i)},\dots,x_{t-1}^{(i)})$ denotes the probability that the model predicts the ground truth label $y_{t}^{(i)}$ at position $t$ in the $i$-th data sample and $T_{i}$ represents the sequence length of the $i$-th data sample.
2.5 Unified Framework and Training
To ensure consistency across the query detection, correction, and answer generation processes, the output of each stage serves as the input for the subsequent stage, ultimately forming a seamless pipeline. Through this three-stage training strategy, the model effectively learns to detect and correct misleading information in queries and generate accurate responses. This approach significantly mitigates the impact and potential harm of misleading information on the model’s performance.
3 Experiments
Instruction Model Method HaluEval
Sum Dia QA HaluEval-QAmis
Qwen2-7B Original $51.5\pm 0.8$ $70.3\pm 0.4$ $54.5\pm 1.3$ $51.1\pm 0.6$
SFT $54.2\pm 1.8$ $71.9\pm 0.5$ $62.3\pm 1.1$ $58.6\pm 1.3$
Ours 68.1 $\pm$ 1.7 72.1 $\pm$ 1.3 68.9 $\pm$ 1.6 68.0 $\pm$ 1.0
Qwen2.5-14B Original $73.2\pm 0.6$ $70.5\pm 0.9$ $62.7\pm 1.7$ $54.3\pm 0.8$
SFT $73.8\pm 0.5$ $71.3\pm 0.6$ $68.2\pm 1.3$ $60.7\pm 1.0$
Ours 76.3 $\pm$ 0.8 72.2 $\pm$ 0.4 73.2 $\pm$ 1.5 71.3 $\pm$ 1.6
Llama3.2-3B Original $27.2\pm 2.3$ $49.9\pm 0.5$ $49.5\pm 1.2$ $46.2\pm 0.7$
SFT $35.0\pm 1.6$ $53.2\pm 1.3$ $51.9\pm 1.0$ $48.1\pm 0.4$
Ours 57.4 $\pm$ 2.3 60.8 $\pm$ 2.1 53.9 $\pm$ 0.8 52.5 $\pm$ 0.9
Llama3-8B Original $50.0\pm 1.8$ $57.3\pm 1.4$ $51.8\pm 0.9$ $49.6\pm 0.6$
SFT $53.8\pm 0.8$ $59.1\pm 1.3$ $57.3\pm 1.3$ $53.2\pm 1.0$
Ours 62.7 $\pm$ 1.5 68.6 $\pm$ 0.6 67.7 $\pm$ 1.8 66.9 $\pm$ 1.4
Table 1: Performance comparison on HaluEval dataset. In this case, Original refers to the model without any processing, while supervised fine-tuning (SFT) refers to the model that undergoes direct fine-tuning.
3.1 Setup
We conducted comprehensive experiments on the hallucination detection task (HaluEval), question answering tasks (CQA and TruthfulQA), and the misleading query datasets we constructed (HaluEval-QAmis and CQAmis). In all experiments, we report the mean results and standard deviations over three random runs. The experimental configuration and dataset usage details are in Appendix A.
Original Datasets
We conduct the construction of the misleading datasets and the evaluation of the model on three datasets:
  • The HaluEval dataset (Li et al., 2023) includes the QA, Summarization, and Dialogue subsets, is used to test the model’s ability to identify hallucinations and misleading information.
  • The CQA dataset (Talmor et al., 2019) is used to evaluate the model’s accuracy in answering common sense questions.
  • The TruthfulQA dataset (Lin et al., 2022) is used to assess the truthfulness and hallucination in the generated answers.
Construction of Training Datasets
Our training dataset consists of 4,000 samples from the QA subset of HaluEval and the corresponding 4,000 samples from the HaluEval-QAmis dataset, which contains misleading information. This dataset is designed to provide diverse training samples that include both genuine and misleading queries. The specific construction process is in Appendix B.
Instruction Model Method CQA CQAmis TruthfulQA
ACC ACC BLEURT BLEU ROUGE
Qwen2-7B Original 70.4 $\pm$ 0.8 68.8 $\pm$ 0.8 0.636 $\pm$ 0.013 0.543 $\pm$ 0.006 0.570 $\pm$ 0.004
SFT 72.1 $\pm$ 0.7 69.9 $\pm$ 1.1 0.633 $\pm$ 0.011 0.562 $\pm$ 0.002 0.577 $\pm$ 0.007
Ours 78.8 $\pm$ 1.6 74.9 $\pm$ 0.5 0.678 $\pm$ 0.006 0.551 $\pm$ 0.003 0.589 $\pm$ 0.006
Qwen2.5-14B Original 80.9 $\pm$ 0.8 71.6 $\pm$ 0.9 0.709 $\pm$ 0.019 0.582 $\pm$ 0.004 0.609 $\pm$ 0.009
SFT 81.4 $\pm$ 0.5 73.3 $\pm$ 1.6 0.721 $\pm$ 0.005 0.577 $\pm$ 0.008 0.613 $\pm$ 0.007
Ours 83.1 $\pm$ 0.7 78.6 $\pm$ 1.3 0.748 $\pm$ 0.008 0.615 $\pm$ 0.011 0.642 $\pm$ 0.013
Llama3.2-3B Original 65.1 $\pm$ 0.9 62.1 $\pm$ 0.6 0.613 $\pm$ 0.006 0.517 $\pm$ 0.005 0.527 $\pm$ 0.005
SFT 66.2 $\pm$ 1.0 63.7 $\pm$ 0.7 0.617 $\pm$ 0.005 0.523 $\pm$ 0.004 0.521 $\pm$ 0.008
Ours 68.5 $\pm$ 0.4 67.5 $\pm$ 0.5 0.632 $\pm$ 0.008 0.534 $\pm$ 0.007 0.556 $\pm$ 0.007
Llama3-8B Original 70.8 $\pm$ 1.1 64.9 $\pm$ 0.5 0.625 $\pm$ 0.011 0.489 $\pm$ 0.012 0.553 $\pm$ 0.004
SFT 69.3 $\pm$ 0.5 64.3 $\pm$ 0.4 0.607 $\pm$ 0.013 0.503 $\pm$ 0.005 0.562 $\pm$ 0.008
Ours 76.4 $\pm$ 1.3 70.0 $\pm$ 0.8 0.668 $\pm$ 0.009 0.528 $\pm$ 0.007 0.586 $\pm$ 0.006
Table 2: Performance comparison on CQA, CQAmis and TruthfulQA dataset.
3.2 Main Results
We conducted experiments on models from the Qwen (Yang et al., 2024) and Llama (Dubey et al., 2024) families. The experimental results for the hallucination detection task are listed in Table 1, and the experimental results for the question answering task are listed in Table 2.
The Impact of Misleading Information in the Input on the Response
For all the original models (those not trained with our framework), the accuracy on the HaluEval-QAmis and CQAmis datasets is significantly lower than that on the QA and CQA datasets. This demonstrates that the presence of misleading information in the input can significantly affect the accuracy and performance of the model’s responses. After SFT, the model shows some improvement in performance on both the original dataset and the misleading query datasets. However, the accuracy on the misleading query datasets HaluEval-QAmis and CQAmis remains significantly lower than that on the original QA and CQA datasets. However, after training with our method, the model’s performance on the misleading query datasets shows a significant reduction in the gap compared to the QA and CQA datasets. This indicates that our proposed method effectively mitigates the performance degradation caused by misleading information in the input.
Performance on the Hallucination Detection Task
We tested the model’s performance on the hallucination detection task using the QA, Summarization, and Dialogue subsets of the HaluEval dataset, as well as the HaluEval-QAmis dataset. The experimental results show that the model trained by our method significantly outperforms the original model as well as the model after SFT on all datasets. This demonstrates that our method enhances the model’s ability to detect misleading information.
Performance on the Question Answering Task
We evaluated the model on the commonsense question answering task using the CQA dataset and the CQAmis dataset. We found that the model trained with our method showed a significant improvement in accuracy on both datasets, demonstrating that our method can enhance the model’s accuracy in question answering tasks. We tested the model’s ability to generate truthful responses and avoid producing false information using the TruthfulQA dataset. The experimental results show that the model trained with our method achieves improvements in BLEURT(Sellam et al., 2020), BLEU (Papineni et al., 2002), and ROUGE (Lin, 2004) metrics, demonstrating that our method can effectively reduce hallucinations in responses and improve the factual accuracy of the model’s answers.
Scalability Analysis
We selected the 7B and 14B models from the Qwen series, as well as the 3B and 8B models from the Llama series, for experimentation. The experimental results show that, regardless of the model family or size, the performance of the models trained with our method is significantly improved. This indicates that our method is applicable to models of different scales and demonstrates good scalability. Furthermore, on the QA, HaluEval-QAmis, and Dialogue datasets, the Qwen-7B model trained with our framework even outperforms the Qwen-14B model that was not trained with our method, further validating the effectiveness and scalability of our approach.
Cross-task Generalisability Analysis
During the model training process, our training data only consisted of 4,000 samples from the QA dataset and 4,000 samples from the HaluEval-QAmis dataset. However, the model still shows significant performance improvements on other datasets. In comparison, we performed SFT on the model using the same training data. However, the model fine-tuned with standard methods showed significant performance improvement only on the QA and HaluEval-QAmis datasets, with little to no improvement on other datasets and even a decline in performance. This demonstrates that our method has strong cross-task generalization capabilities.
3.3 Instruction Following Capability Study
We evaluated the model’s instruction-following capability on the three subtasks of the hallucination detection dataset HaluEval and the HaluEval-QAmis dataset. In these evaluations, if the model’s response does not include either “YES” or “NO”, or contains both, it is marked as failed. Table 3 shows the proportion of failed responses for each model in this evaluation task. The models trained with our method exhibit a significant reduction in the failed proportion across all datasets, indicating that our method significantly improves the model’s instruction-following capability.
Instruction Model Method HaluEval
Sum Dia QA HaluEval-QAmis
Llama3.2-3B Original 13.67 1.96 3.54 5.67
SFT 8.73 2.17 2.36 2.86
Ours 3.55 1.24 1.09 2.09
Llama3-8B Original 2.61 0.63 1.03 1.75
SFT 2.17 0.58 0.95 1.42
Ours 0.36 0.43 0.69 0.78
Table 3: Performance comparison of failure rate (%) on HaluEval dataset. A smaller failure rate (%) indicates a stronger ability of LLMs to follow instructions.
Instruction Model Method HaluEval CQA
Sum Dia QA HaluEval-QAmis CQA CQAmis
Llama-3.2-3B Original $27.2\pm 2.3$ $49.9\pm 0.5$ $49.5\pm 1.2$ $46.2\pm 0.7$ $65.1\pm 0.9$ $62.1\pm 0.6$
Stage 1 $52.8\pm 1.0$ $56.4\pm 1.5$ $50.1\pm 1.0$ $48.4\pm 0.6$ $67.6\pm 0.6$ $66.1\pm 0.9$
Stage 2 $53.2\pm 0.7$ $57.2\pm 1.3$ $50.5\pm 0.7$ $48.7\pm 0.4$ $68.1\pm 0.8$ $66.5\pm 0.4$
Stage 3 57.4 $\pm$ 2.3 60.8 $\pm$ 2.1 53.9 $\pm$ 0.8 52.5 $\pm$ 0.9 68.5 $\pm$ 0.4 67.5 $\pm$ 0.5
Llama3-8B Original $50.0\pm 1.8$ $57.3\pm 1.4$ $51.8\pm 0.9$ $49.6\pm 0.6$ $70.8\pm 1.1$ $64.9\pm 0.5$
Stage 1 $60.1\pm 0.6$ $65.3\pm 1.2$ $63.1\pm 1.2$ $62.1\pm 1.2$ $73.5\pm 1.2$ $67.4\pm 1.1$
Stage 2 $60.5\pm 0.8$ $65.7\pm 0.8$ $63.6\pm 0.9$ $63.0\pm 0.7$ $74.8\pm 1.9$ $68.9\pm 0.7$
Stage 3 62.7 $\pm$ 1.5 68.6 $\pm$ 0.6 67.7 $\pm$ 1.8 66.9 $\pm$ 1.4 76.4 $\pm$ 1.3 70.0 $\pm$ 0.8
Table 4: Ablation study on HaluEval and CQA dataset. Original refers to the original model, Stage 1 refers to the model after query detection training, Stage 2 refers to the model after further query correction training, and Stage 3 refers to the model after further query answering training again.
3.4 Ablation Study
We used the 3B and 8B versions of the Llama models to test the changes in model performance after each stage of training. The results are shown in Table 4.
The Role of the Query Detection Stage
As can be seen from the results in Table 4, the models after query detection training show significant performance improvement on all datasets. Moreover, the larger the model size, the more significant the performance improvement. This verifies that the query detection stage effectively enhances the model’s ability to match query details with existing knowledge, thereby enabling it to identify potential misleading information more accurately. Therefore, after fine-tuning with the query detection stage, the model’s ability to identify misleading information is significantly enhanced, providing a solid foundation for subsequent hallucination correction and answer generation.
The Role of the Query Correction Stage
As can be seen from the results in Table 4, the models, after further query correction training, continue to exhibit performance improvements across all datasets. In Appendix D, we present a case study that illustrates that after this step of training, the model has learned to correct queries containing misleading information.
The Role of the Answer Generation Stage
As can be seen from the results in Table 4, the models after further query answering training show significant improvement both on the hallucination detection task and the question answering task. The main role of the answer generation training is to help the model adapt to the corrected queries, enabling it to better understand the modified query and generate more accurate answers that align with the user’s true intention. Therefore, this stage not only greatly improves the model’s performance but also further consolidates the improvements achieved through the previous two fine-tuning stages, making the model more accurate when handling complex tasks.
3.5 Case Study
We provide case studies as examples in Table 16 and Appendix D, which visually demonstrate that the model trained with our method can correctly identify and correct the misleading information in the question, compared to the original model, thereby generating the correct answer. This significantly improves the accuracy of the responses and reduces hallucinations.
In the example from the HaluEval-QAmis dataset (Table 16), when faced with a query containing misleading information, the original model responds directly based on the misleading question, resulting in an incorrect answer “FA Premier League”. Although this competition is indeed a British football event, it does not align with the competition that the question actually intends to ask about, which is the one attended by Malcolm Smith. This indicates that misleading information can mislead the model into providing an incorrect response. In contrast, the model trained with our method keenly detects the misleading information in the question, corrects the misleading information, and then generates a response based on the corrected question, yielding the correct result.
HaluEval-QAmis Example       Ground Truth Label: Super Bowl XLVIII
Query "knowledge": " Smith was named the Most Valuable Player of Super Bowl XLVIII after they defeated the Denver Broncos. Super Bowl XLVIII was an American football game between the American Football Conference (AFC) champion Denver Broncos and National Football Conference (NFC) champion Seattle Seahawks to decide the National Football League (NFL) champion for the 2013 season." "question": "In which England football game was Malcolm Smith named Most Valuable player?"
Model Output
(Original model)
FA Premier League
Model Output
(Our Method)
Malcolm Smith named Most Valuable Player is American football game, not England football game, this American football game is Super Bowl XLVIII.
Table 5: Case study on the HaluEval-QAmis dataset. The original model is misled by the misleading information "England football" in the query, resulting in an incorrect answer (highlighted in red). However, the model trained with our method correctly identifies that it should be “American football” and generates the correct answer (highlighted in blue).
4 Related Work
The Impact of Misleading Information on LLMs
Studies have shown that in question answering systems, when the input contains false or ambiguous information, LLMs often generate inaccurate responses based on these erroneous inputs (Pan et al., 2023; Zhang et al., 2024b; Kuhn et al., 2022; Chen and Shu, 2024). To address this issue, a common approach is to detect potential errors or inconsistencies based on the syntactic and semantic analysis of the input. For example, Li and Wang (2024) proposed a detection-correction integrated framework based on a general language model. This approach integrates the detection and correction processes into a single model, enabling the automatic detection and correction of syntactic errors in the input text. To ensure the accuracy of the input information, some studies (Zhang et al., 2024a; Butala et al., 2024; Xiong et al., 2024) have attempted to validate the correctness of the input through multi-turn dialogues, ensuring that the content generated by the model aligns with reality. Although existing methods have improved the model’s ability to handle misleading information to some extent, the complexity and diversity of misleading information still pose significant challenges in this area.
Model Output Correction
In recent years, methods aimed at correcting the content generated by LLMs to improve generation accuracy have been widely studied. Wu et al. (2024b) proposed an iterative verification-correction framework (ProCo), which identifies and corrects potential errors through subsequent validation of the generated content. Petroni et al. (2021) leveraged external knowledge bases to correct the content generated by the model. Liu et al. (2024a) explored the intrinsic self-correction capabilities of LLMs from both theoretical and empirical perspectives, proposing that zero temperature and fair prompting are key factors for successful self-correction. Liu et al. (2024b) proposed distilling the LLMs’ self-evaluation capability and more comprehensive thinking into smaller models, effectively improving the performance and answer accuracy of LLMs in resource-constrained environments. In addition, knowledge-enhanced generation has also been used to reduce the likelihood of language models generating misleading content (Jiang et al., 2024).
5 Conclusion
In this paper, we propose a novel three-stage fine-tuning method to mitigate the impact of misleading information in the input queries on LLMs. Our method enhances the ability of LLMs to detect, correct, and respond accurately to queries containing misleading content, significantly improving response accuracy and reducing hallucinations. Extensive experiments across various datasets demonstrate the effectiveness of this approach, making LLMs more robust and trustworthy.
6 Limitations
Although our method has achieved significant improvement in mitigating the impact of misleading information in the input on LLMs, there are still some limitations that can be addressed in future work:
  1. 1.
    Dependency on high-quality knowledge: Our method relies on the model’s internal knowledge or requires reliable external knowledge during the query detection and Query Correction Stages. If both the model’s internal knowledge and the external knowledge base are incomplete or inaccurate, the model may still generate incorrect corrections or responses.
  2. 2.
    Scalability to larger models: Although our method performs well across different model sizes (e.g., Qwen-7B vs. Qwen-14B, Llama 3B vs. 8B models), future work can explore experiments on even larger models.
  3. 3.
    Expansion of the training dataset: Future work could consider expanding the training dataset to more effectively improve model performance and facilitate transfer to additional tasks or domains.
7 Ethical Statement
The Use of AI Assistants
We employed ChatGPT to assist us in polishing our paper and coding.
Appendix A Experimental details
We employ LoRA for efficient fine-tuning. The detailed setting of hyperparameters is shown in Table 6. Table 7 presents the information about the datasets used for training and testing. Regarding SFT, we use the QA and HaluEval-QAmis datasets for implementation.
Configuration Value
Number of epochs 4
Devices 4 Nvidia H20 GPU (80G)
Total Batch size 128
Learning rate $5\times 10^{-4}$
Warmup Ratio 0.1
Table 6: Finetuning hyperparameters for experiments.
Datasets Train number Test number
QA 4000 6000
HaluEval-QAmis 4000 6000
Summarization 0 10000
Dialogue 0 10000
CQA 0 1221
CQAmis 0 1221
TruthfulQA 0 818
Table 7: Dataset Split.
Appendix B Dataset details
The specific process of constructing our training dataset is as follows:
  • Query Detection Stage: We used a mixed dataset comprising QA and HaluEval-QAmis samples for training. Additionally, we introduced a field $with-misleading$ as a label to indicate whether a query contains misleading information. For queries in the QA dataset, the $with\text{-}misleading$ field is set to “NO”, indicating that the query does not contain misleading information. For queries in the HaluEval-QAmis dataset, the $with\text{-}misleading$ field is set to “YES”, indicating that the query contains misleading information.
  • Query Correction Stage: Since the task in this stage requires correcting queries that contain misleading information, we selected 4,000 samples from the HaluEval-QAmis dataset for training. At this point, our training objective is to correct the misleading queries and transform them into the corresponding queries from the original QA dataset.
  • Answer Generation Stage: In this stage, we also use a mixed dataset of QA and HaluEval-QAmis samples to train the model to generate correct answers, where the queries containing misleading information have already been corrected in the query correction stage.
This dataset construction strategy ensures that the model can focus on the corresponding data characteristics at each task stage, while also enhancing the model’s ability to handle diverse queries. Next, we present the prompts that we used in the experiments as well as the HaluEval-QAmis and CQAmis datasets. Table 8 and Table 10 show the templates used to generate the HaluEval-QAmis and CQAmis datasets, while Table 9 and Table 11 display some data from the HaluEval-QAmis and CQAmis datasets. Table 12, Table 13, and Table 14 respectively show the templates for query detection, misleading query correction, and answer generation.
Prompt for Generating HaluEval-QAmis Dataset
You are a language expert. You should create a question that includes subtle mistakes derived from the content in the knowledge. The modified question should retain the general meaning of the original but introduce small errors or misdirecting details such as incorrect entity names, relationships, times, or terminology. These changes should be subtle and should not drastically alter the original question’s structure or core intent. Do not introduce completely new concepts, terms, or laws that may cause the mistake question to deviate too far from the original. The mistake question should still be recognizable as a version of the original question, with enough misleading elements to potentially cause an incorrect answer if not identified and corrected. The goal is to create a subtle version of the original question that misleads the model, but does not make it so different that the answer is hard to recover. Do not provide any content, only the misleading question. """ #knowledge#: {knowledge} #question#: {question} """
Table 8: The prompt for generating the HaluEval-QAmis dataset.
The example of HaluEval-QAmis dataset
Example1: "knowledge": "The 2003 SEC Championship Game was won by the LSU Tigers 34–13 over the Georgia Bulldogs. The game was played in the Georgia Dome in Atlanta, Georgia. Located in Atlanta, Georgia, between downtown to the east and Vine City to the west, it was owned and operated by the State of Georgia as part of the Georgia World Congress Center Authority." "question": "The 2003 SEC Championship Game was played at a stadium owned and operated by who?" "misleading question": "The 2003 SEC Championship Game was played at a stadium in the United States owned and operated by who?" "right answer": "the State of Georgia as part of the Georgia World Congress Center Authority" "hallucinated answer": "The film P̈andorumẅas released in 2010." "The 2003 SEC Championship Game was played at the Mercedes-Benz Stadium in Atlanta."
Example2: "knowledge": " He was also a successful author, writing biographies of politician Al Smith and athletes Lou Gehrig and John McGraw, as well as histories of the New York Yankees, New York Giants and Brooklyn Dodgers. The team’s name derived from the reputed skill of Brooklyn residents at evading the city’s trolley streetcar network." "question": "What team’s name derived from the reputed skill of Brooklyn residents at evading the city’s trolley streetcar network, who had a history wrote of it by an American sportswriter and biographer who overed sports in New York for the N̈ew York Sunf̈rom 1915 to 1943?" "misleading question": "What team’s name derived from the reputed skill of Manhattan residents at evading the city’s trolley streetcar network?" "right answer": "Brooklyn Dodgers" "hallucinated answer": "The New York Yankees were named after the reputed skill of Brooklyn residents at evading the city’s trolley streetcar network."
Example3: "knowledge": "Sooni Taraporevala (born 1957) is an Indian screenwriter, photographer and filmmaker who is best known as the screenwriter of M̈ississippi Masala,̈ T̈he Namesakeänd Oscar-nominated S̈alaam Bombay(̈1988), all directed by Mira Nair. Mississippi Masala is a 1991 romantic drama film directed by Mira Nair, based upon a screenplay by Sooni Taraporevala, starring Denzel Washington, Sarita Choudhury, and Roshan Seth." "question": "Which Oscar-nominated film was written by the screenwriter of a 1990 romantic drama starring Sarita Choudhury?" "right answer": "Salaam Bombay" "hallucinated answer": "The Namesake"
Table 9: Randomly sampled HaluEval-QAmis dataset examples. Blue represents correct information, while red represents misleading information.
Prompt for generating CQAmis dataset
You are a language expert. Your task is to take a given question and modify it by introducing distracting or irrelevant details to make the problem more complex and harder to answer. These distractions should not directly lead to any of the answer choices, especially the correct one. Steps: 1. Modify the Question: Add irrelevant or distracting details that make the question more challenging. These distractions can be excessive descriptions, background information, or other elements that do not directly relate to the core of the question. 2. Provide a Detailed Explanation: After modifying the question, provide a clear explanation that shows why the correct answer is still valid. The explanation should demonstrate that the added distractions don’t affect the reasoning behind the correct answer and why it is still the best choice. Important Notes: The distractions should increase the complexity of the question, but should not directly or indirectly guide the reader to the correct answer. The explanation should clearly outline why the correct answer is the best choice and how the added distractions do not change the fundamental reasoning. Example: """ "question": "A thoroughfare meandered through fields and woods, where was it passing though?", "choices": "["move about","city","country","town","new york city"]". """ "mistake question": "This road winds through various landscapes, sometimes open farmlands, other times dense forests, and occasionally small villages. Along the way, there are signs of modern construction like small factories and even a few billboards. The surroundings shift frequently, with the occasional sight of tall buildings in the distance. Where is this road most likely passing through?" "explanation": "The road is passing through country because, despite the modern construction and occasional tall buildings in the distance, the key elements—farmlands, forests, and small villages are characteristic of rural, countryside areas. The mention of "small factories" and "billboards" are distractions that suggest a more urban environment, but they do not outweigh the overall rural characteristics of the area described." """ #question#: {question} #choices#: {choices} """
Table 10: The prompt for generating the CQAmis dataset. In addition to the original fields, we also generated an explain field to provide a detailed description of how the question, after adding misleading information, is step by step deduced to the correct answer.
The example of CQAmis dataset
Example1: "question": "Billy was reading the newspaper as he commuted to work, but once he got to his destination he balled it up and put it somewhere. Where did it put it?" "choices": ["trash", "floor", "subway", "ground", "lawn"] "answer": "trash" "misleading question": Billy was engrossed in reading the morning newspaper during his daily commute. As he arrived at his workplace, he noticed the crowded train and decided to quickly stuff the newspaper into his pocket to avoid carrying it around. However, as the day progressed, he realized he had forgotten about it and didn’t know where he had placed it. Where could it be now?" "explanation": Despite the added details of Billy stuffing the newspaper into his pocket and forgetting about it later, the core question remains about where Billy put the newspaper once he reached his destination. The hallucinations do not provide any information about the newspaper’s final location; thus, the correct answer, ẗrash,r̈emains valid. Billy might have thrown away the newspaper if he didn’t need it anymore or if he found it cluttering his belongings."
Example2: "question": "He had to wear a tuxedo while playing the keyboard instrument, so did the other hundred members of the what?" "choices": ["music store", "band", "medium", "orchestra", "piano store"] "answer": "orchestra" "misleading question": "He had to wear a tuxedo while playing the keyboard instrument, and the other hundred members also wore tuxedos. There was a piano beside him. Which group did they belong to?" "explanation": "Although there was a piano beside him, the fact that they had to wear tuxedos still leads to the correct answer being "orchestra."
Table 11: Randomly sampled CQAmis dataset examples. Red represents misleading information.
Prompt for question detection
You are an excellent mistake detector. I will provide you with a #knowledge# and a #question#. Your task is to carefully compare the #knowledge# and the #question#, and based on the content of the #knowledge#, determine whether the #question# contains any mistake. If it contains a mistake, output ’YES’, otherwise output ’NO’. Please make sure there are no extra outputs. The following is an example: #knowledge#: "The area covered by South East Queensland varies, depending on the definition of the region, though it tends to include Queensland’s three largest cities: the capital city Brisbane; the Gold Coast; and the Sunshine Coast. The Gold Coast is a coastal area in the Australian state of Queensland, approximately 66 km south-southeast of the state capital Brisbane and immediately north of the border with New South Wales.". #question#: "South East Queensland (SEQ) is a bio-geographical, political, and administrative region of the state of Queensland in Australia, the area covered by South East Queensland varies, depending on the definition of the region, though it tends to include Queensland’s three largest cities, including which coastal area in the Australian state of Victoria, approximately 66 km south-southeast of the state capital Melbourne and immediately north of the border with New South Wales?". #your choice#:"YES" You need to do your best to identify whether there is any content in the #question# that contradicts the #knowledge#, and if there is, output "YES", otherwise output "NO".
Table 12: The prompt for question detection.
Prompt for misleading question correction
You are an excellent mistake corrector. I will provide you with a #knowledge# and a #question#. Your task is to carefully compare the #knowledge# and the #question#, and based on the content of the #knowledge# and real-world information, correct any mistakes in the #question# and output the corrected content. The following is an example: #knowledge#: "The area covered by South East Queensland varies, depending on the definition of the region, though it tends to include Queensland’s three largest cities: the capital city Brisbane; the Gold Coast; and the Sunshine Coast. The Gold Coast is a coastal area in the Australian state of Queensland, approximately 66 km south-southeast of the state capital Brisbane and immediately north of the border with New South Wales." #question#: "South East Queensland (SEQ) is a bio-geographical, political, and administrative region of the state of Queensland in Australia, the area covered by South East Queensland varies, depending on the definition of the region, though it tends to include Queensland’s three largest cities, including which coastal area in the Australian state of Victoria, approximately 66 km south-southeast of the state capital Melbourne and immediately north of the border with New South Wales?". #your answer#: "South East Queensland (SEQ) is a bio-geographical, political, and administrative region of the state of Queensland in Australia, the area covered by South East Queensland varies, depending on the definition of the region, though it tends to include Queensland’s three largest cities, including which coastal area in the Australian state of Queensland, approximately 66 km south-southeast of the state capital Brisbane and immediately north of the border with New South Wales?" You need to carefully identify if there is any content in the #question# that contradicts the #knowledge# or real-world knowledge. If there is, please modify the #question# to the correct form and output it. Otherwise, directly output the original #question#.
Table 13: The prompt for misleading question correction.
Prompt for generate answer
You are an expert with broad knowledge. I will provide #knowledge# and #question#. Your task is to answer the #question# based on the #knowledge# and real-world information, and output the answer to the #question#. The following is an example: #knowledge#: "He was one of the pioneers of gambling in Las Vegas, where he was a partner with flamboyant mobster Bugsy Siegel at the Flamingo Hotel. Benjamin B̈ugsyS̈iegel (February 28, 1906 – June 20, 1947) was a Jewish American mobster. Siegel was not only influential within the Jewish mob but, like his friend and fellow gangster Meyer Lansky, he also held significant influence within the Italian-American Mafia and the largely Italian-Jewish National Crime Syndicate.". #question#: "What hotel did the influential National Crime Syndicate member found with David Berman?". #your answer#:"Flamingo Hotel" You need to ensure that the answer you provide contains no content that contradicts the #knowledge# or real-world knowledge."
Table 14: The prompt for generating an answer.
Appendix C Select misleading question
This section demonstrates how to select high-quality misleading questions using sentence similarity and answer error rate, with the detailed process shown in Table 15.
HaluEval-QAmis Example       Ground Truth Label:1965
Knowledge DuPont v. Kolon Industries is an intellectual property lawsuit centering on the allegation that Kolon Industries (of ), a South Korea-based company, stole trade secrets concerning the production and marketing of Kevlar from DuPont, an American chemical company. Kevlar is a high strength synthetic fiber used in applications as diverse as bicycle tires and body armor. Developed by Stephanie Kwolek at DuPont in 1965, this high-strength material was first commercially used in the early 1970s as a replacement for steel in racing tires.
Original Query What year was the high strength synthetic fiber, which was the subject of a lawsuit between DuPont and Kolon Industries, developed?
Misleading Query 1 What year was the low strength synthetic fiber, which was the subject of a lawsuit between DuPont and Kolon Industries, developed?
Similarity Score:0.96
respond1: 1867 ()
respond2: 1965 ()
respond3: 1433 ()
Error Rate:0.67
Misleading Query 2 What year was the high strength synthetic fiber, which was the subject of a lawsuit between DuPont and Japan Industries, developed?
Similarity Score:0.96
respond1: 1965 ()
respond2: 1965 ()
respond3: 1965 ()
Error Rate:0
Misleading Query 3 What year was the high strength synthetic fiber, which was the subject of a lawsuit between China and England Industries, developed?
Similarity Score:0.92
respond1: 1965 ()
respond2: 1965 ()
respond3: 1563 ()
Error Rate:0.33
Table 15: An example of randomly generated misleading queries and the calculation of $S_{\text{sim}}$ and $E_{\text{error}}$, where the misleading information is highlighted in red. Example 1 satisfies our filtering condition of $S_{\text{sim}}$ > 0.8 and $E_{\text{error}}$ > 0.5. Correct answers are marked with a correct mark , while incorrect answers are marked with a wrong mark . Among them, misleading query 1 satisfies our filtering criteria of $S_{\text{sim}}$ > 0.8 and $E_{\text{error}}$ > 0.5.
Appendix D Case study
Similar to the analysis in Section 3.5, in the example of the CQAmis dataset (Table 17), the query mentions that after the mechanic made some adjustments to the engine, the car started to experience slight vibrations and emit some unusual sounds. These distracting pieces of information can easily lead the model to mistakenly interpret the sounds as coming from the horn, resulting in an incorrect answer “honk the horn”. However, the model trained with our method effectively filters out the interference from the misleading information and provides a response stating that the ‘unusual sounds’ and ‘slight vibrations’ are simply distractions that might suggest the car is malfunctioning, but the fact that it’s operational and no serious issues are mentioned indicates the car will start running as expected after repairs.
We present a case study of the query correction stage in Table 18, the model trained using our method is able to quickly detect misleading information in the question. By leveraging the knowledge from “the last man executed in Australia”, it correctly modifies the misleading information in the query, “the last person executed in New Zealand”, and generates the correct question.
HaluEval-QAmis Example       Ground Truth Label: Super Bowl XLVIII
Query "knowledge": " Smith was named the Most Valuable Player of Super Bowl XLVIII after they defeated the Denver Broncos. Super Bowl XLVIII was an American football game between the American Football Conference (AFC) champion Denver Broncos and National Football Conference (NFC) champion Seattle Seahawks to decide the National Football League (NFL) champion for the 2013 season." "question": "In which England football game was Malcolm Smith named Most Valuable player?"
Model Output
(Original model)
FA Premier League
Model Output
(Our Method)
Malcolm Smith named Most Valuable Player is American football game, not England football game, this American football game is Super Bowl XLVIII.
Table 16: A case study on the HaluEval-QAmis dataset. The original model is misled by the misleading information "England football" in the question, resulting in an incorrect answer (highlighted in red). However, the model trained with our method correctly identifies that it should be “American football” and generates the correct answer (highlighted in blue).
CQAmis Example       Ground Truth Label: start running
Query "question": "After the mechanic performed several adjustments to the engine, including checking the fuel system and replacing some critical parts, the car started to behave differently. It began to vibrate slightly and made some unusual sounds. Although it seems operational, it’s not the same as it was before. What is it most likely to do next?" "choices": "["going too fast","last several years","honk the horn","go fast","start running"]"
Model Output
(Original model)
honk the horn
Model Output
(Our Method)
Even though the car is vibrating and making unusual sounds, these are typical signs of a car being started after maintenance, especially if adjustments were made to the engine. The "unusual sounds" and "slight vibrations" are simply distractions that might suggest the car is malfunctioning, but the fact that it’s operational and no serious issues are mentioned indicates the car will start running as expected after repairs.
Table 17: A case study on the CQAmis dataset. The original model is misled by the misleading information "slight vibrations" and "unusual sounds" in the question, resulting in an incorrect answer (highlighted in red). However, the model trained with our method correctly identifies "it’s operational and no serious issues are mentioned" and generates the correct answer "the car will start running" (highlighted in blue).
Category Content
Knowledge Input His most well-known work is the award-winning stage play ’Remember Ronald Ryan’, a dramatization of the life and subsequent death of Ronald Ryan, the last man executed in Australia. Ronald Joseph Ryan (21 February 1925 – 3 February 1967) was the last person to be legally executed in Australia.
Original Question Barry Dickins wrote a film about the last person executed in New Zealand, named what?
Output The knowledge mentions the last person legally executed in Australia, not New Zealand, so the question should be modified to "Barry Dickins wrote a film about the last person executed in Australia, named what?"
Table 18: An example of correcting misleading inputs based on knowledge. For the input data, we used the same prompt as in the query correction training stage (Table 13). The information marked in blue represents correct information, while the information marked in red represents misleading information.